OpenStudy (anonymous):

Use the result of the previous problem to find the Laplace Transform of the given function:

OpenStudy (anonymous):

\[\large f(t)=\left\{\begin{matrix} 1, & 0 \le t <1\\ 0, & 1 \le t <2 \end{matrix}\right.\]\[\large f(t+2)=f(t)\] `PREVIOUS PROBLEM:` Show that for f(t+T)=f(t):\[\large \mathcal{L}[f(t)]=\frac{\int\limits_{0}^{T} e^{-st}f(t)dt}{1-e^{-sT}}\]

OpenStudy (anonymous):

@dan815 @amistre64 @Directrix

OpenStudy (irishboy123):

my bandwidth is so bad that i cannot see this properly, but i think you need the transform of the starting function before you go to the second bit. guy to ask is @ganeshie8

OpenStudy (anonymous):

Hey @IrishBoy123 ! Let me get a screenshot! And yeah but he's not online at the moment! =/

OpenStudy (anonymous):

Do you see why \(T=2\)?

OpenStudy (anonymous):

With that in mind, \[\mathcal{L}\{f(t)\}=\frac{\displaystyle\int_0^2e^{-st}f(t)\,\mathrm{d}t}{1-e^{-2s}}=\frac{\displaystyle\int_0^1e^{-st}\,\mathrm{d}t}{1-e^{-2s}}=\cdots\]

OpenStudy (anonymous):

@SithsAndGiggles I'm not sure I understand what you did there =/

OpenStudy (anonymous):

Oh I think I see. How did you get rid of f(t) though?

OpenStudy (anonymous):

Just taking advantage of the definition of \(f(t)\). Over the interval \((0,1)\), you have \(f(t)=1\), while over \((1,2)\), \(f(t)=0\). You can thus split up the integral like so: \[\int_0^2e^{-st}f(t)\,\mathrm{d}t=\int_0^1e^{-st}\times1\,\mathrm{d}t+\int_1^2e^{-st}\times0\,\mathrm{d}t=\int_0^1e^{-st}\,\mathrm{d}t\]

OpenStudy (anonymous):

Ahhh yeah that makes sense. Thanks!!